Prove that for every positive integer β€˜n’ : 1𝑋2𝑋3 + 2𝑋3𝑋4 + β‹― + 𝑛(𝑛 + 1)(𝑛 + 2) =


𝑛(𝑛+1)(𝑛+2)(𝑛+3)


1
Expert's answer
2022-05-06T05:42:49-0400

Let "P(n)" be the proposition thatΒ 


"1\times 2\times 3+2\times 3\times 4+...+n(n+1)(n+2)"

"=\dfrac{n(n+1)(n+2)(n+3)}{4}"

Basis step: "P(1)" is true, because


"1\times 2\times 3=6=\dfrac{1(1+1)(1+2)(1+3)}{4}"

Inductive step: For the inductive hypothesis we assume that "P(k)" holds for an arbitrary positive integer "k." That is, we assume that


"1\times 2\times 3+2\times 3\times 4+...+k(k+1)(k+2)"

"=\dfrac{k(k+1)(k+2)(k+3)}{4}"

Under this assumption, it must be shown that "P(k + 1)" is true, namely, that


"1\times 2\times 3+2\times 3\times 4+...+k(k+1)(k+2)"

"+(k+1)(k+1+1)(k+1+2)"


"=\dfrac{(k+1)(k+1+1)(k+1+2)(k+1+3)}{4}"

is also true.

We obtain


"1\times 2\times 3+2\times 3\times 4+...+k(k+1)(k+2)"

"+(k+1)(k+1+1)(k+1+2)"


"=\dfrac{k(k+1)(k+2)(k+3)}{4}+(k+1)(k+2)(k+3)"

"=\dfrac{(k+1)(k+2)(k+3)}{4}\cdot(k+4)"

"=\dfrac{(k+1)(k+1+1)(k+1+2)(k+1+3)}{4}"


This last equation shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all positive integers n. That is, we have proven that


"1\times 2\times 3+2\times 3\times 4+...+k(k+1)(k+2)"

"=\dfrac{k(k+1)(k+2)(k+3)}{4}"

for all positive integers "n." Β